A ball tossed h feet in the air will stay in the air for T= 0: 5p (square root then) h

seconds. Note that when a question uses units the final answer should include appropriate units.


(a) Suppose a ball is tossed 4 feet in the air. When will the ball come down?


(b) If a ball remains in the air for 2.5 seconds, how high was it tossed


1
Expert's answer
2021-09-20T08:39:55-0400

"T=\sqrt{h}, h" in "ft" , "t" in "sec"

(a) Given "h=4 \ ft"

Then

"T=\sqrt{4} \ sec=2\ sec"

(b) Given "T=2.5\ sec"

Then

"h=(2.5)^2\ ft=6.25\ ft"